# Motion blur; Blind deblurring; Blind motion deblurring; Deep learning

### v1

@@ Abstract
Motion deblurring is one of the fundamental problems of computer vision and has received continuous attention. The variability in blur, both within and across images, imposes limitations on non-blind deblurring techniques that rely on estimating the blur kernel. As a response, blind motion deblurring has emerged, aiming to restore clear and detailed images without prior knowledge of the blur type, fueled by the advancements in deep learning methodologies. Despite strides in this field, a comprehensive synthesis of recent progress in deep learning-based blind motion deblurring is notably absent. This paper fills that gap by providing an exhaustive overview of the role of deep learning in blind motion deblurring, encompassing datasets, evaluation metrics, and methods developed over the last six years. Specifically, we first introduce the types of motion blur and the fundamental principles of deblurring. Next, we outline the shortcomings of traditional non-blind deblurring algorithms, emphasizing the advantages of employing deep learning techniques for deblurring tasks. Following this, we categorize and summarize existing blind motion deblurring methods based on different backbone networks, including convolutional neural networks, generative adversarial networks, recurrent neural networks, and Transformer networks. Subsequently, we elaborate not only on the fundamental principles of these different categories but also provide a comprehensive summary and comparison of their advantages and limitations. Qualitative and quantitative experimental results conducted on four widely used datasets further compare the performance of state-of-the-art methods. Finally, an analysis of present challenges and future pathways aims to guide and inspire researchers in the dynamic domain of image deblurring using deep learning techniques. This review aspires to serve as an invaluable reference for those immersed in this field of study. All collected models, benchmark datasets, source code links, and codes for evaluation have been made publicly available at https://github.com/VisionVerse/Blind-Motion-Deblurring-Survey
@@
Motion blur; Blind deblurring; Blind motion deblurring; Deep learning
[…]
@@ 1 Introduction
Motion blur is a common artifact in images caused by the relative movement between the camera and the scene during the exposure Burdziakowski (2020); Peng et al. (2021). This effect can cause object contours in the image to be blurred or stretched, thereby reducing the clarity and detail of the image and affecting various computer vision tasks such as autonomous driving Rengarajan et al. (2016); Li et al. (2023), object segmentation Rajagopalan et al. (2023); Zhou et al. (2023) and scene analysis Zhang and Bergholm (1997); Zhou et al. (2022). Understanding the origins of motion blur is critical to developing effective elimination strategies.
The blurry process, as a fundamental degradation model, can be depicted as the convolution between a clear image and a blur kernel Chen et al. (2021); Bai et al. (2019); Zhang et al. (2022):
@@
where $\mathbf{B}$ and $\mathbf{S}$ are the observed blurred image and sharp image, respectively. $\mathbf{k}$ and $\mathbf{n}$ represent the blur kernel and noise. $\otimes$ is the convolution operation. When $\mathbf{k}$ is known, the inverse process of this task is referred to as non-blind deblurring Arjomand Bigdeli et al. (2017); Sun et al. (2014); Zhang et al. (2017). However, $\mathbf{k}$ is often unknown when dealing with blurred images, leading to tasks known as blind deblurring Pan et al. (2016); Chen et al. (2019); Hu et al. (2014); Yan et al. (2017); Liu et al. (2019).
Motion blur can be divided into two types: global motion blur and local motion blur. As shown in Figure 1 (b), global motion blur is caused by the shaking of the camera Zhai et al. (2023), resulting in consistent blurring across the entire image. As shown in Figure 1 (c), local motion blur occurs due to relative motion between the target object and the static background, resulting in blur Zhang et al. (2022).
Figure 1: Sharp-blurry images.
@@
Deblurring represents an ill-posed problem that cannot be directly solved. The traditional blind deblurring methods Levin (2006); Yang and Ji (2019); Fergus et al. (2006) start with the blur degradation model, employing mathematical modeling or convolutional networks to predict the blur kernel. Then they utilize the estimated blur kernel to restore the image through deconvolution. These approaches rely on prior knowledge of both the image and the blur kernel to address the ill-posedness in deblurring. As shown in the traditional method on Figure 2, the commonly used prior algorithms are sparsity prior Krishnan et al. (2011); Xu et al. (2013); Levin et al. (2011), patch prior Sun et al. (2013); Guo and Ma (2018); Tang et al. (2018), channel prior Pan et al. (2017); Ge et al. (2021); Cai et al. (2020), gradient prior Levin et al. (2009); Pan et al. (2014); Shao et al. (2020) and so on. The quality of image deblurring is ensured by the priori algorithms. These traditional methods heavily rely on blur kernel estimation, and the effectiveness of the deblurring process directly depends on the accuracy of the kernel estimation. However, in complex real-world scenarios, the blur kernel in the blurring process tends to be complex, variable, and difficult to parameterize, causing the traditional methods to often perform inadequately.
In recent years, deep learning technology has developed rapidly and has been widely applied in image processing Zhang et al. (2012); Zhou et al. (2024). Deep learning-based methods have powerful feature learning and mapping capabilities, allowing them to learn complex patterns of blur removal from extensive datasets Gu et al. (2020). Consequently, significant progress has been made in image deblurringZhao et al. (2022); Purohit et al. (2021); Chen et al. (2022); Yang and Yamac (2022); Nimisha et al. (2017) As depicted in Figure 3, existing blind motion deblurring methods can be categorized based on different backbone networks, namely: convolutional neural networks (CNN) Mao et al. (2016), generative adversarial networks (GAN) Goodfellow et al. (2014); Wang et al. (2023), recurrent neural networks (RNN) Zhang et al. (2018), and Transformer networks Zhao et al. (2023). The deep learning methods can adaptively learn the blur features according to the training data, thus achieving end-to-end image deblurring. They can directly generate sharp images from blurred ones, which improves the effect of image deblurring. Compared to the traditional methods, the deep learning methods Li et al. (2021); Yuan et al. (2020); Noroozi et al. (2017) demonstrate improved robustness in real-world scenarios and possess broader applicability.
@@
  • In this review, we comprehensively overview the research progress of deep learning in blind motion deblurring, en-compassing the causes of motion blur, blurred image datasets, evaluation metrics for image quality, and methods developed over the past six years.
@@
  • Furthermore, for each type of blind motion deblurring method, we provide an in-depth summary of their fundamental principles, analyzing the advantages and limitations of each approach. Qualitative and quantitative experimental results on four widely used datasets further showcase the performance differences among different blind motion deblurring methods.
  • Finally, we conduct a prospective analysis and summary of current research challenges and potential future directions in blind motion deblurring. This review aims to offer researchers and practitioners in the field a comprehensive under-standing and a valuable reference for further advancements.
Figure 3: Overview of deep learning methods for blind motion deblurring.
The subsequent sections of this paper are structured as follows: We first summarized the classification and research status of image blind motion deblurring methods in Sec. 2. Then, We briefly introduced the public datasets and evaluation metrics in Sec. 3. Next, Sec. 4 describes the basic principles of different types of methods in detail, and summarizes and compares the advancements and shortcomings of different methods. Qualitative and quantitative experimental results on four common datasets further compare the performance of different methods. Finally, in Sec. 5 and Sec. 6, we provide a prospective analysis of current research challenges and potential future directions in blind motion deblurring.
@@ 2.1 CNN-based Blind Motion Deblurring Methods
CNN-based image blind deblurring methods can be divided into two categories according to the process of image deblurring. One is the early two-stage image deblurring network shown in Figure 4 (a), and the other is the end-to-end image deblurring methods with the better performance shown in Figure 4 (b).
@@ 2.1.1 Two-stage CNN-based Blind Motion Deblurring Methods
Classical methods such as: Schuler et al. Schuler et al. (2015) trained a blind image deconvolution network to deblur the image. This network primarily consists of three parts: a feature extraction module, a blur kernel estimation module, and an image estimation module. The estimated blur kernel from the blur kernel estimation module is used to iteratively update the deblurred image. Sun et al. Sun et al. (2015) proposed to estimate the heterogeneous motion blur through the CNN model and then deconvolute the blurred images, which can effectively estimate the spatially varying motion kernel and achieve a better effect of removing motion blur. To address the issue of the non-uniform blur, Cronje et al. Cronje (2015) proposed a CNN-based image deblurring method. This method trained two CNN networks, one for estimating the vector of non-uniform motion and the other for image deconvolution. It accurately estimated non-uniform blur and enhanced the effectiveness of image deblurring. Ayan Chakrabarti et al. Chakrabarti (2016) trained a neural network for estimating the blur kernel of blur images. Using the estimated single global blur kernel, they used non-blind deconvolution to obtain deblurred images. Gong et al. Gong et al. (2017) trained a fully convolutional neural network for estimating image motion flow. Then they preformed non-blind deconvolution to recover clear images from the estimated motion flow. This method avoids the iterative process of underlying image priors. Xu et al. Xu et al. (2017) proposed a deep convolutional neural network to extract sharp edges from blur images for kernel estimation, which does not require any coarse-to-fine strategy or edge selection, thus greatly simplifying the kernel estimation and reducing the computational amount. Kaufman et al. Kaufman and Fattal (2020) analyze the blurring kernel by applying convolutional layers and pooling operations to extract features at multiple scales. Subsequently, the estimated kernel, along with the input image, is fed into a synthetic network to perform non-blind deblurring.
These two-stage image deblurring methods rely too much on the blur kernel estimation of the first stage, and the estimated blur kernel quality is directly related to the effect of image deblurring. In reality, the blur is mostly uneven, and the size and direction of the blurred image are also uncertain. So in the real scenes, this method can not remove complex real blur very well.
@@
2.1.2 End-to-end CNN-based Blind Motion Deblurring Methods
The end-to-end image deblurring method generates a clear image directly from the input blurred image. It utilizes neural networks to learn complex feature mapping relationships for better restoration of clear images. End-to-end image deblurring algorithms have made significant progress since Xu et al. Xu et al. (2014) first implemented end-to-end motion blur image restoration using CNNs. As shown in Figure 5, CNN-based end-to-end image deblurring methods can be categorized into three types: multi-scale, multi-patch, and multi-temporal structures.
@@ 2.1.2 End-to-end CNN-based Blind Motion Deblurring Methods
Multi-scale Structure: As shown in Figure 5 (a), the multi-scale model extracts multi-scale feature information from blurry images. Based on this extracted information, it progressively restores the image from coarse to fine levels. In order to better handle non-uniform blur in dynamic scenes, Nah et al. Nah et al. (2017)proposed a multi-scale convolutional neural network for end-to-end restoration of blurred images. Following a similar approach, many subsequent researchers have improved on the multi-scale strategy to obtain better results Purohit and Rajagopalan (2020); Tao et al. (2018); Vitoria et al. (2022); Hemanth and Latha (2020). However, its limitation lies in scaling the blurred image to a lower resolution, often resulting in the loss of edge information. Overlaying multi-scale input images onto sub-networks gradually enhances image clarity from lower-level to higher-level sub-networks, inevitably incurring higher computational costs. To reduce the spatio-temporal complexity of employing a multi-scale strategy network, Cho et al. Cho et al. (2021) devised a network architecture that starts from coarse to fine, proposing a multi-input multi-output U-Net network. This approach allows for a better execution of deblurring from coarse to fine.
@@
Multi-patch Structure: As shown in Figure 5 (b), the multi-patch structure segments the blurred input into multiple patches for prediction. Inspired by spatial pyramid matching, Zhang et al. Zhang et al. (2019) proposed a deep hierarchical end-to-end CNN model called DMPHN. This model used multiple patches for a fine-to-coarse operation to handle blurry images. To remove spatially varying blur, Suin et al. Suin et al. (2020) proposed a patch hierarchical attention architecture. It significantly enhances deblurring effects through effective pixel-wise adaptation and feature attention. Building upon DMPHN, Zhang et al. Zhang et al. (2023) proposed a self-supervised event-guided deep hierarchical multi-patch network, which handles blurry images through a layered representation from fine to coarse. By stacking patch networks to increase network depth, this model enhanced deblurring performance. However, when segmenting input or features, discontinuities in contextual information may arise, potentially generating artifacts and leading to suboptimal image deblurring in dynamic scene contexts Zamir et al. (2021).
@@
Multi-temporal Structure: As shown in Figure 5 (c), the multi-temporal structure gradually eliminates non-uniform blur over multiple iterations. The multi-temporal framework Park et al. (2020) was introduced to handle non-uniform blur and exhibited significant performance improvements. However, rigid progressive training and inference processes might not adapt well to images with varying degrees of blur in different regions. Subsequent researchers have proposed numerous time-series-based deblurring methods, learning from different temporal scales of motion blur to achieve enhanced deblurring effects Zhang et al. (2023a, b).

### v2

# Deep Learning in Motion Deblurring: Current Status, Benchmarks and Future Prospects

@@ Abstract
Motion deblurring is one of the fundamental problems of computer vision and has received continuous attention. The variability in blur, both within and across images, imposes limitations on non-blind deblurring techniques that rely on estimating the blur kernel. As a response, blind motion deblurring has emerged, aiming to restore clear and detailed images without prior knowledge of the blur type, fueled by the advancements in deep learning methods. Despite strides in this field, a comprehensive synthesis of recent progress in deep learning-based blind motion deblurring is notably absent. This paper fills that gap by providing an exhaustive overview of the role of deep learning in blind motion deblurring, encompassing datasets, evaluation metrics, and methods developed over the last six years. Specifically, we introduce the types of motion blur and outline the shortcomings of traditional non-blind deblurring algorithms, emphasizing the advantages of employing deep learning techniques for deblurring tasks. Following this, we categorize and summarize existing blind motion deblurring methods based on different backbone networks, including convolutional neural networks, generative adversarial networks, recurrent neural networks, and Transformer networks. Subsequently, we elaborate not only on the fundamental principles of these different categories but also provide a comprehensive summary and comparison of their advantages and limitations. Qualitative and quantitative experimental results conducted on four widely used datasets further compare the performance of state-of-the-art methods. Finally, an analysis of present challenges and future prospects is intended to foster advancements and drive innovation in the field of image deblurring research. This review aspires to serve as a reference for those immersed in this field of study. All collected models, benchmark datasets, source code links, and codes for evaluation have been made publicly available at https://github.com/VisionVerse/Blind-Motion-Deblurring-Survey.
@@
Motion blur, Blind deblurring, Blind motion deblurring, Deep learning
@@ 1 Introduction
Motion blur is a common artifact in images caused by the relative movement between the camera and the scene during the exposure [1, 2]. This effect can cause object contours in the image to be blurred or stretched, thereby reducing the clarity and detail of the image. It not only affects the visual effect of images, but also poses a challenge to a series of computer vision tasks such as autonomous driving [3, 4], object segmentation [5, 6], medical image processing[7, 8] and scene analysis [9]. To address this challenge, we need to construct a model that can describe the process of motion blur, thus providing us with a framework for understanding the nature of motion blur.
Figure 1: Sharp-blurry images.
The blurry process, as a fundamental degradation model, can be depicted as the convolution between a clear image and a blur kernel [10, 11, 12]:
@@
where $\mathbf{B}$ and $\mathbf{S}$ are the observed blurred image and sharp image, respectively. $\mathbf{k}$ and $\mathbf{n}$ represent the blur kernel and noise. $\otimes$ is the convolution operation. When blur kernel $\mathbf{k}$ is known, the inverse process of this task is referred to as non-blind deblurring [13, 14, 15, 16, 17, 18, 19]. However, $\mathbf{k}$ is often unknown when dealing with blurred images, leading to tasks known as blind deblurring [20, 21, 22, 23, 24, 25, 26, 27]. Motion blur can be divided into two types: global motion blur and local motion blur. As shown in Figure 1 (b), global motion blur is caused by the shaking of the camera [28], resulting in consistent blurring across the entire image. As shown in Figure 1 (c), local motion blur occurs due to relative motion between the target object and the static background [29].
Deblurring represents an ill-posed problem that cannot be directly solved. The traditional blind deblurring methods [30, 31, 32, 33] start with the blur degradation model, employing mathematical modeling or convolutional networks to predict the blur kernel. Then they utilize the estimated blur kernel to restore the image through deconvolution. These approaches rely on prior knowledge of both the image and the blur kernel to address the ill-posedness in deblurring. As shown in the traditional method on Figure 2, the commonly used prior algorithms are sparsity prior [34, 35, 36], patch prior [37, 38, 39], channel prior [40, 41, 42], gradient prior [43, 44, 45] and so on. These traditional methods heavily rely on blur kernel estimation, and the effectiveness of the deblurring process directly depends on the accuracy of the kernel estimation. However, in complex real-world scenarios, the blur kernel in the blurring process tends to be complex, variable, and difficult to parameterize, causing the traditional methods to often perform inadequately [16]. Deblurring methods based on deep learning can learn the nonlinear mapping of blurred images to clear images end-to-end. The effect of inaccurate blur kernel estimation on the deblurring is avoided. Deep learning methods tend to have better deblurring results in dealing with non-uniform blurring problems in the real world.
@@
In recent years, deep learning technology has developed rapidly and has been widely applied in image processing [46, 47, 48, 49, 50, 51, 52]. Deep learning-based methods have powerful feature learning and mapping capabilities, allowing them to learn complex patterns of blur removal from extensive datasets [53]. Consequently, significant progress has been made in image deblurring[54, 55, 56, 57, 58].
As depicted in Figure 3, existing blind motion deblurring methods can be categorized based on different backbone networks, namely: convolutional neural networks (CNN) [59, 60, 61], generative adversarial networks (GAN) [62, 63], recurrent neural networks (RNN) [64], and Transformer networks [65]. The deep learning methods can adaptively learn the blur features according to the training data. They can directly generate sharp images from blurred ones, which improves the effect of image deblurring. Compared to the traditional methods, the deep learning methods [66, 67, 68] demonstrate improved robustness in real-world scenarios and possess broader applicability.
@@
  • In this review, we comprehensively overview the research progress of deep learning in blind motion deblurring, encompassing the causes of motion blur, blurred image datasets, evaluation metrics for image quality, and methods developed over the past six years.
@@
  • For each type of blind motion deblurring method, we provide an in-depth summary of their fundamental principles, analyzing the advantages and limitations of each approach. Qualitative and quantitative experimental results on four widely used datasets further showcase the performance differences among different blind motion deblurring methods.
  • Finally, we conduct a prospective analysis and summary of current research challenges and potential future directions in blind motion deblurring. This review aims to offer researchers and practitioners in the field a comprehensive understanding and a valuable reference for further advancements.
Figure 3: An overview of deep learning methods for blind motion deblurring.
The subsequent sections of this paper are structured as follows: We first summarized the classification and research status of image blind motion deblurring methods in Sec. 2. Then, We briefly introduced the public datasets and evaluation metrics in Sec. 3. Next, Sec. 4 describes the basic principles of different types of methods in detail, and summarizes and compares the advancements and shortcomings of different methods. Qualitative and quantitative experimental results on four common datasets further compare the performance of different methods. Subsequently, in Sec. 5, we provide a prospective analysis of current research challenges and potential future directions in blind motion deblurring. Finally, in Sec. 6, we provide a summary of our paper.
@@ 2.1 CNN-based Blind Motion Deblurring Methods
CNN-based image blind deblurring methods can be divided into two categories according to the process of image deblurring. One is the early two-stage image deblurring framework shown in Figure 4 (a), and the other is the end-to-end image deblurring framework with the better performance shown in Figure 4 (b).
@@ 2.1.1 Two-stage CNN-based Blind Motion Deblurring Methods
Schuler et al. [69] trained a blind image deconvolution network to deblur the image. This network primarily consists of three parts: a feature extraction module, a blur kernel estimation module, and an image estimation module. The estimated blur kernel from the blur kernel estimation module is used to iteratively update the deblurred image. Sun et al. [70] proposed to estimate the heterogeneous motion blur through the CNN model and then deconvolute the blurred images, which can effectively estimate the spatially varying motion kernel and achieve a better effect of removing motion blur. To address the issue of the non-uniform blur, Cronje et al. [71] proposed a CNN-based image deblurring method. This method trained two CNN networks, one for estimating the vector of non-uniform motion and the other for image deconvolution. It accurately estimated non-uniform blur and enhanced the effectiveness of image deblurring. Ayan Chakrabarti et al. [72] trained a neural network for estimating the blur kernel of blur images. Using the estimated single global blur kernel, they performed deconvolution operation on the blur images to obtain deblurred images. Gong et al. [73] trained a fully convolutional neural network for estimating image motion flow. Then they preformed non-blind deconvolution to recover clear images from the estimated motion flow. This method avoids the iterative process of underlying image priors. Xu et al. [74] proposed a deep convolutional neural network to extract sharp edges from blur images for kernel estimation, which does not require any coarse-to-fine strategy or edge selection, thus greatly simplifying the kernel estimation and reducing the computational amount. Kaufman et al. [75] analyze the blurring kernel by applying convolutional layers and pooling operations to extract features at multiple scales. Subsequently, the estimated kernel, along with the input image, is fed into a synthetic network to perform non-blind deblurring.
These two-stage image deblurring methods rely too much on the blur kernel estimation of the first stage, and the estimated blur kernel quality is directly related to the effect of image deblurring. In reality, the blur is mostly uneven, and the size and direction of the blur are also uncertain. So in the real scenes, this method can not remove complex real blur very well.
2.1.2 End-to-end CNN-based Blind Motion Deblurring Methods
The end-to-end image deblurring method generates a clear image directly from the input blurred image. It utilizes neural networks to learn complex feature mapping relationships for better restoration of clear images [76]. End-to-end image deblurring algorithms have made significant progress since Xu et al. [77] first implemented end-to-end motion blur image restoration using CNNs. As shown in Figure 5, CNN-based end-to-end image deblurring methods can be categorized into three types: multi-scale, multi-patch, and multi-temporal structures.
@@ 2.1.2 End-to-end CNN-based Blind Motion Deblurring Methods
Multi-scale Structure:As shown in Figure 5 (a), the multi-scale model extracts multi-scale feature information from blurry images. Based on this extracted information, it progressively restores the image from coarse to fine levels[78]. In order to better handle non-uniform blur in dynamic scenes, Nah et al. [79]proposed a multi-scale convolutional neural network for end-to-end restoration of blurred images. Following a similar approach, many subsequent researchers have improved on the multi-scale strategy to obtain better results [80, 81, 82, 83]. However, its limitation lies in scaling the blurred image to a lower resolution, often resulting in the loss of edge information. Overlaying multi-scale input images onto sub-networks gradually enhances image clarity from lower-level to higher-level sub-networks, inevitably incurring higher computational costs. Dong et al. [84] proposed a simple and effective multi-scale residual low-pass filtering network, which can adaptively explore the spatially varying properties of the global context and effectively model both high-frequency and low-frequency information. To reduce the spatio-temporal complexity of employing a multi-scale strategy network, Cho et al. [85] devised a network architecture that starts from coarse to fine, proposing a multi-input multi-output U-Net network. This approach allows for a better execution of deblurring from coarse to fine. To address the oversight of image frequency distinction based on multi-scale methods, Zhang et al. [86] proposed a multi-scale frequency separation network. By combining multi-scale strategies with frequency separation, it enhances the restoration of details in blurry images.
@@
Multi-patch Structure: As shown in Figure 5 (b), the multi-patch structure segments the blurred input into multiple patches for prediction. Inspired by spatial pyramid matching, Zhang et al. [87] proposed a deep hierarchical end-to-end CNN model called DMPHN. This model used multiple patches for a fine-to-coarse operation to handle blurry images. To remove spatially varying blur, Suin et al. [88] proposed a patch hierarchical attention architecture. It significantly enhances deblurring effects through effective pixel-wise adaptation and feature attention. Building upon DMPHN, Zhang et al. [89] proposed a self-supervised event-guided deep hierarchical multi-patch network, which handles blurry images through a layered representation from fine to coarse. By stacking patch networks to increase network depth, this model enhanced deblurring performance. However, when segmenting input or features, discontinuities in contextual information may arise, potentially generating artifacts and leading to suboptimal image deblurring in dynamic scene contexts [90].
@@
Multi-temporal Structure: As shown in Figure 5 (c), the multi-temporal structure gradually eliminates non-uniform blur over multiple iterations. The multi-temporal framework [91] was introduced to handle non-uniform blur and exhibited significant performance improvements. However, rigid progressive training and inference processes might not adapt well to images with varying degrees of blur in different regions. Subsequent researchers have proposed numerous time-series-based deblurring methods, learning from different temporal scales of motion blur to achieve enhanced deblurring effects [92, 93].
2.2 RNN-based Blind Motion Deblurring Methods
Recurrent Neural Networks (RNNs) are commonly used for handling sequential data such as text, language, and time series, demonstrating exceptional performance when dealing with data that exhibit temporal or sequential patterns [94]. Due to their memory-based neural network architecture, RNNs have found applications in image processing as well. For instance, Zhang et al. [64] proposed an end-to-end spatially variant RNN network for deblurring in dynamic scenes. As illustrated in Figure 6, the weights of RNNs are learned by deep CNNs. They analyzed the relationship between spatially variant RNNs and the deconvolution process, demonstrating the ability of spatially variant RNNs to simulate the deblurring process. Models trained using the proposed RNNs are notably smaller and faster. To better utilize the properties of spatially varying RNNs, Ren et al. [95] extend a unidirectionally connected one-dimensional spatially varying RNN to a two-dimensional spatial RNN with three-way connectivity. 2D RNNs can capture a larger receptive field and learn denser propagation between neighboring pixels compared to 1D RNNs, further improving deblurring performance. Tao et al. [81] proposed a multi-scale deep neural network for single image deblurring based on convolutional long short-term memory (Conv LSTM), effectively utilizing previous frames and image features. Dongwon Park et al. [91] developed a multi-temporal recurrent neural network (MT-RNN) based on recursive feature mapping for blind single-image deblurring, progressively deblurring a single image in an iterative process. Zhang et al. [84] proposed an effective dynamic scene deblurring method based on optical flow and spatial variable RNN, which utilizes optical flow to guide the spatially variable RNN to remove blur. In practical applications, CNNs and other models are often combined with RNNs to compensate for certain limitations in image processing, such as sequence length restrictions and computational complexity.